OpenStudy (pinknabastak):

HELP!!!!! A plane moves at a certain velocity and then accelerates at the rate of 52m/s^2 for a distance of 2300m to reach a velocity of 620m/s. How fast was the plane going before it began to accelerate?

OpenStudy (javk):

are you studying mechanics?

OpenStudy (pinknabastak):

its for my physics class

OpenStudy (paxpolaris):

do you have an equation that doesn't use time?

OpenStudy (anonymous):

here you need to find the initial velocity

OpenStudy (pinknabastak):

uhh yea...its vf^2=vi^2+2ad

OpenStudy (javk):

ok so u need to use the eqn: \[v ^{2}= u ^{2}+2as\]

OpenStudy (anonymous):

yup

OpenStudy (pinknabastak):

so 620=vi^2+2(52)(2300)

OpenStudy (pinknabastak):

i mean 620^2

OpenStudy (pinknabastak):

so 38440=vi^2+239200

OpenStudy (pinknabastak):

so then square root of 145200 which is 381.05 m/s

OpenStudy (pinknabastak):

which is my answer right?

OpenStudy (javk):

yup

OpenStudy (pinknabastak):

THANKS! :)